Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2017-09-30

Observers: Kris Rybicki, Simon Hodgkin
TAT: Regis Cartier, Janet Chen
Support team: Morgan Fraser, Emma Callis

Calibrations: biases, dome arcs+flats

Conditions: Clear overhead - cloud looks like its all on the horizon. Seeing 1.5ish arcsec

Problem identified in the afternoon with the telescope Azimuth. Not fixed by sunset.
23:36 UT - telescope fixed, and setup beginning (for SOFI)
00:20 Clear outside, seeing not great
01:50 Refocus
02:20 Looks clear out
Sunrise: clear



UT       Target            Type            Comment
==============================================================================================     
00:21    AT2017gah        JHK Pho
02:05    SN2016iks        JH Pho
03:34    SN2016iks        K Pho
04:31    ASASSN-15og      JH Pho
06:00    ASASSN-15og      K Pho
07:00    AT2017fzw        JHK Pho
08:20    AT2017fzw        GR11+GR16 FUP
09:24    GD71             STD

Comments